### v2

# supplement

Abstract
Predictive simulation of surface chemistry is of paramount importance for progress in fields from catalysis to electrochemistry and clean energy generation. Ab-initio quantum many-body methods should be offering deep insights into these systems at the electronic level, but are limited in their efficacy by their steep computational cost. In this work, we build upon state-of-the-art correlated wavefunctions to reliably converge to the ‘gold standard’ accuracy in quantum chemistry for application to extended surface chemistry. Efficiently harnessing graphics processing unit acceleration along with systematically improvable multiscale resolution techniques, we achieve linear computational scaling up to 392 atoms in size. These large-scale simulations demonstrate the importance of converging to these extended system sizes, achieving a validating handshake between simulations with different boundary conditions for the interaction of water on a graphene surface. We provide a new benchmark for this water-graphene interaction that clarifies the preference for water orientations at the graphene interface. This is extended to the adsorption of carbonaceous molecules on chemically complex surfaces, including metal oxides and metal-organic frameworks, where we consistently achieve chemical accuracy compared to experimental references, and well inside the scatter of traditional density functional material modeling approaches. This pushes the state of the art for simulation of molecular adsorption on surfaces, and marks progress into a post-density functional era for more reliable and improvable approaches to first-principles modeling of surface problems at an unprecedented scale and accuracy using ab-initio quantum many-body methods.
1 Main
Figure 1: Systematically improvable quantum embedding (SIE) framework and our GPU-accelerated quantum embedding for surfaces. Panel (a) illustrates the general workflow for a controllable multi-scale resolution of correlations at different length scales, comprising SIE+MP2 (purple box), SIE+CCSD (orange box), and SIE+CCSD(T) (red box), with further details in the Methods Sec. 4 and supplemental information (SI) Sec. LABEL:sec:SI_method. Panel (b) presents the computational time as a function of the total number of orbitals, measured on a single A100 GPU for systems comprising water monomer adsorbed on graphene under open boundary condition. The annotations beside the curves show the scaling obtained by fitting the last three points up to the maximum size of the test system. Further discussion on the observed linear scaling of SIE+CCSD(T) are provided in SI section LABEL:sec:SI_LS_SIE.
The chemical and physical properties of material surfaces play a central role in a wide range of scientific research areas and industrial applications [1, 2, 3]. However, predictive simulations of these systems without resorting to empiricism remains a formidable and long-standing challenge, requiring an accurate solution to the many-electron problem [4, 5, 6, 7, 8, 9]. The formal solution to this problem scales exponentially in computational effort with system size, and therefore to make progress necessitates a sufficiently accurate approximation to the electron correlations while also rendering the problem computationally efficient to scale to large systems at the bulk limit. Over the past few decades, density functional theory (DFT) has emerged as the standard bearer for this, favored for its wide applicability and low computational cost [10, 11, 12]. However, DFT is not systematically improvable due to its reliance on semi-empirical exchange-correlation functionals, that are not universal and may not provide sufficient transferability or internal validation of its accuracy across different chemical environments. In contrast, correlated methods [13], such as coupled cluster theories [9] or quantum Monte Carlo methods [8], can achieve superior accuracy by explicitly describing electron correlation, and importantly their accuracy can be systematically improved [14, 15, 16, 17, 18, 19]. Among these, the coupled-cluster with single, double, and perturbative triple excitations (CCSD(T)) [20], is often regarded as the ‘gold standard’ of electronic structure theory, and has been shown to accurately model the relevant long-range interactions on material surfaces [21, 22, 23]. However, in the worst case, the correlation present within the system could span across hundreds of atoms, and this poses a significant challenge for CCSD(T) due to its steep computational scaling with both system and basis set size, severely limiting its applicability for realistic surface chemistry.
This work considers advances that address the challenges of simulating realistic material surfaces with these quantum many-body methods, as well as using these developments to validate the accuracy and enable key insights into a chemically diverse range of outstanding surface absorption problems. This advance in simulation capabilities leverages the growing field of quantum embedding to introduce a controllable locality approximation that achieves a practical linear scaling in computational effort for system sizes considered [24]. Specifically, we extend the previously proposed ‘systematically improvable quantum embedding’ (SIE) method [25, 26], which itself builds on density matrix embedding theory and fragmentation methods of quantum chemistry [27, 28, 21, 29, 30, 31]. Our developments extend the approach to couple together layers of different resolutions of correlated effects at different length scales, up to the ‘gold standard’ CCSD(T) level. We demonstrate exceptional accuracy with linear scaling, crucial for performing these simulations at the required scale. We also adapt the approach to utilize graphics processing units (GPUs) to eliminate key computational bottlenecks in our workflow, including the implementation of GPU-enhanced correlated solvers. With these, we can converge unprecedented simulations at the CCSD(T) level over solid-state systems of tens of thousands of orbitals, as summarized by the workflow of Fig. 1a and the observed computational scaling of Fig. 1b, with more details of these methodological advances described in Sec. 4 and Supporting Information (SI) Sec. LABEL:sec:SI_method.
2 Results
2.1 Water molecule on graphene
Figure 2: The adsorption energy of water on graphene at the bulk limit. The computed interaction energies ($\Delta E_{\text{int}}$) of H2O@Graphene are shown for both 0-leg (green) and 2-leg (purple) configurations (shown top right) as well as with open (circles) and periodic (triangles) boundary conditions. These are obtained for different system sizes with both the SIE+CCSD(T) workflow, and compared to other methods. The infinity symbol represents an extrapolation to the bulk limit. The final absorption energies, $\Delta E_{\text{ads}}$, shown in the inset table combine both an extrapolation of the computed interaction energies to the bulk limit at the SIE+CCSD(T) level, with a geometry relaxation correction at DFT level. The superscript a represents results taken from Ref. [32].
The adsorption of water on surfaces plays a central role in fields from desalination [33] to clean energy [34, 35] and many more besides. The case of graphene represents a remarkably fundamental surface for this process, with its semi-metallic native state harboring a number of novel properties and promising applications including blue energy [35] and quantum friction [36]. However, the dominating weak and long-range van der Waals interactions between water and graphene lead to a highly non-local interaction that poses significant technical challenges to achieve convergence with respect to the size of the graphene sheet [37, 38, 39, 40, 32, 41, 42]. A method capable of reliably computing the adsorption energies of water on graphene (H2O@Graphene) therefore serves as a meaningful indicator of the feasibility of simulating realistic surface chemistry more broadly.
This limiting finite-size error can be qualitatively estimated via the difference between adsorption energies calculated for structures of similar size under both open and periodic boundary conditions (OBC and PBC respectively), regarded as the OBC-PBC gap [32]. The OBC-PBC gap arises because the OBC and PBC models exhibit a different convergence of their finite-size errors, stemming from their contrasting physical origins. In OBC models, the error arises from the artificially truncated range of the interactions of the finite-sized substrate with the adsorbate, whereas under PBC conditions it is caused by the spurious periodic interactions between all the particles with their images in neighbouring cells. Our developments provide an opportunity to confidently eliminate the finite-size error via a handshake between the absorption energies computed under these different boundary conditions.
[…]
Figure 3: Adsorption of water on graphene at different orientations. (a) Interaction energy of water adsorption on graphene at various orientation angles $\theta$. Calculations are performed at the SIE+CCSD level under OBC using PAH structures ranging from PAH(2) to PAH(8). (b) The cross-sectional views of the electron density rearrangement for PAH(6) with different water orientations. (c) Top views of the isosurface of electron density rearrangement, plotted with a cutoff value of $6.0\times 10^{-5}$ electrons/Å3.
The 0- and 2-leg configurations are unique cases in which the dipole moment of the water molecule is perpendicular to the graphene surface. Water adsorption on graphene with varying orientations has not been extensively explored, but represents an important problem to address to understand the dynamics of water across graphene and subsequent implications in tribology applications. We can characterize the orientation of the water by an angle of rotation, $\theta$, with the 2-leg configuration defined at $\theta=0^{\circ}$ and the 0-leg configuration defined at $\theta=180^{\circ}$, with additional computational details in the SI Sec. LABEL:sec:SI_distance_opt_water_orientation. SIE+CCSD calculations were conducted on PAH(h) under OBC with $h=2,4,6,8$, and the results are shown in Fig. 3a. Again, we observe significant finite-size effects which change both the relative ordering and absolute scales of these absorption energies over the different orientations. In particular, we find that the long-range interaction depends significantly on the orientation of water and is maximized for the 0-leg and 2-leg configurations. Furthermore, long-range interactions stabilize the adsorption for $\theta>60^{\circ}$, and destabilizes the adsorption for $\theta<60^{\circ}$. This emphasizes the importance of being able to converge to the bulk limit for all orientations to correctly describe their relative stability. Notably, for the configuration at $\theta=60^{\circ}$ (red line), the interaction energy remains nearly constant as the graphene substrate size increases, indicating that finite-size errors are particularly small in this specific case.
For more insight into these observations, we analyze the electron density rearrangement (see the definition in SI Sec. LABEL:sec:SI_EDR) for the 2-leg, 0-leg, and $\theta=60^{\circ}$ configurations, with their cross-sectional side views shown in Fig. 3b, and isosurface top views in Fig. 3c, respectively. This adsorption-induced electron density rearrangement is considered as another descriptor for interaction range. For the 2-leg configuration, the dipole of the water points toward the substrate, pulling electrons from the graphene to the adsorption site and enhancing the water-graphene interaction, as shown in Fig. 3 b and c. The electron density rearrangement around the absorbed site converges to the bulk limit faster than the free surface, thereby weakening the interaction energy as the system size increases to the converged limit for $\theta<60^{\circ}$. For the 0-leg configuration, electrons in the substrate are pushed away from the adsorption site. Larger substrates allow for the repelled electrons to more effectively rearrange and relax, thereby lowering the energy of the absorbed system and strengthening the interaction towards bulk limits for $\theta>60^{\circ}$. For the special case of $\theta=60^{\circ}$, a balance is reached between these two effects as electrons are pulled from one side and pushed to the other side, as shown in Fig. 3c. This balance largely cancels out the finite-size effects of the interaction energy. However, this does not mean that the effects of the interaction for $\theta=60^{\circ}$ are particularly short ranged. As can be seen from the electron density rearrangement in Fig. 3 b and c, the extent of density changes in the $\theta=60^{\circ}$ system is not smaller than that for the 0-leg and 2-leg, and this cancellation for rapid convergence of the interaction energy with system size is likely to be only found for energetic properties. As a counter-example, we can also compute the adsorption-induced dipole moment (shown in SI Sec. LABEL:sec:SI_DP). Unlike the interaction energy, it is found that the adsorption-induced dipole moment for the $\theta=60^{\circ}$ configuration exhibits pronounced changes as a function of substrate size. Based on this we can conclude that a full description of the interaction-driven effects between water and graphene are all long ranged.
The substantial differences in finite-size convergence for these rotations underline the importance of being able to rigorously converge the size of the graphene substrate before conclusions can be drawn. Considering that the extrapolated SIE+CCSD(T) absorption energies for the 0-leg and 2-leg configurations are already within their associated uncertainties of each other, it is expected that the adsorption energies among all configurations could be comparably close. This means that graphene does not exhibit a strong preference for any specific water orientation, which is unusual for water, as the polar and directional hydrogen-bonding nature of water molecules usually results in a pronounced preference for specific adsorption configurations on surfaces [43, 44]. Such behavior suggests that water-graphene interaction is dominated by non-directional van der Waals forces, unlike the interaction between water and other surfaces [44, 45, 46, 47, 48]. This is supported by weak interaction analysis [49, 50, 51] presented in SI Sec. LABEL:sec:SI_weak_interaction_analysis. Recent studies of quantum friction phenomena  [52, 53, 36] and anomalous dielectric response [54, 55] at water-graphitic interfaces indicate that the standing descriptions of water at these interfaces based on lower-level theories are insufficient. Our results further highlight the necessity to pursue a high-level theoretical description of the water-graphene interface to explain these counter-intuitive emergent phenomena.
2.2 Carbonaceous molecules on various surfaces
Figure 4: Structures and adsorption energies for carbonaceous molecules on different systems. (a) CO on a 4-layer MgO(001) surface with PBC. (b) 6 organic molecules used as the adsorbate on graphene which is replaced by Coronene under OBC. (c) CO/CO2 adsorption configurations in CPO-27-Mg. The 6B cluster is cut from CPO-27-Mg, used to calculate interacting energies under OBC. Structures are rendered by using VESTA [56]. (d) Adsorption energies obtained through SIE+CCSD(T) (stars), while other methods shown with hollow triangles; An inset shows differences between SIE+MP2/CCSD/CCSD(T) and experimental data in kcal/mol. The references are labeled as superscript from (a-j) corresponding to published data from references [57, 58, 47, 59, 60, 61, 62, 63, 64, 65], respectively, while k refers to this work.
Carbonaceous molecules are ubiquitous in nature and synthetic materials. The adsorption of these carbonaceous molecules plays a crucial role in various fields, such as pollution control, drug design and catalyst optimization [66, 67, 68, 69, 70]. As a key quantity in determining surface chemistry, adsorption energies have been computationally probed for various carbonaceous molecules and surfaces to understand their behavior [62, 59, 71, 63, 65, 64, 72, 47, 73, 45, 74, 75, 76, 77, 78, 79]. To demonstrate the wide applicability of our method, we consider three representative systems with chemically diverse surfaces selected: carbon monoxide on the MgO(001) surface, denoted as ‘CO@MgO’, six organic molecules on coronene, a simplified model to mimic graphene, denoted as ‘Organic molecules@Coronene’ [62], and carbon monoxide [63]/carbon dioxide [65, 64] on the metal-organic framework CPO-27-Mg, denoted as ‘CO/CO2@CPO-27-Mg’. Fig. 4a-c shows the adsorption configurations and the surfaces/clusters. The adsorption energies across these systems range from $4$ to $14$ kcal/mol. This range spans from typical weak physisorption to strong chemisorption, highlighting the diversity of these distinct surface types.
To demonstrate the robustness and versatility of the SIE+CCSD(T) protocol, we compare our results to a variety of computational approaches across the systems. Detailed descriptions are provided in the SI Sec. LABEL:sec:SI_benchmark, and we present only the main results here. The computed values, compared to their corresponding experimental data, are shown in Fig. 4d, where the black line represents perfect agreement between theory and experiment, and the gray shade indicates deviation within a ‘chemical accuracy’ of ±$\pm 1$ kcal/mol. We note that the accuracy of the same method, for example MP2, can vary among different adsorption systems. Furthermore, even in the same system (such as CO@MgO), different implementations can result in significant differences, due to the convergence of technical parameters. Notably, the SIE+CCSD(T) method achieves a sub-chemical accuracy in agreement with experimental data across all studied systems.
The inset in Fig. 4d further illustrates the deviations of SIE+MP2, SIE+CCSD, and SIE+CCSD(T) relative to experimental values, highlighting the significant improvement in accuracy of SIE+CCSD(T) over SIE+MP2 and SIE+CCSD. Another important reason is that SIE+CCSD(T) maintains the size consistency inherent in Coupled Cluster theory, which is crucial for achieving consistent high accuracy across systems of different sizes. A contrasting example is MP2, which is widely known as lacking size consistency. Although it performs very well in systems involving the adsorption of small molecules, such as CO/CO2@CPO-27-Mg and CO@MgO, in the case of Organic molecules@Coronene, where all six types of organic molecules are larger than the CO/CO2 examples, MP2 performs poorly. As described in SI Sec. LABEL:sec:SI_OM@Coronene, as the size of organic molecules increases, the differences between MP2 and experimental values significantly widen, whereas the differences between SIE+CCSD(T) and experimental values do not show as marked a change. This consistent high-accuracy performance bolsters our confidence that SIE+CCSD(T) possesses good versatility and transferability to be applied across complex surfaces and with a wide variety of adsorbates.
3 Conclusions and outlook
High-accuracy simulations of realistic material surfaces using quantum many-body methods inevitably encounter two main challenges: (1) accurately capturing the essential electron correlations within materials, and (2) favorable computational scaling to reach the bulk limit, which can involve hundreds of atoms and tens of thousands of basis functions or more to account for long-range interactions. In this work, we demonstrate that our highly efficient GPU-accelerated SIE+CCSD(T) framework effectively addresses these challenges for typical adsorption systems. For water on graphene — a famously challenging system due to its pronounced long-range interactions — we scaled our calculations to converge to CCSD(T)-level accuracy through a systematic computational approach. This allowed us to eliminate finite-size errors which stymied previous studies. At the bulk limit, the adsorption energy converges to approximately 100 meV for both the 0-leg and 2-leg configurations, both lying inside the uncertainty of the calculations. The interaction energies for different orientations of water are also within this range, suggesting for the first time that graphene is insensitive to the orientation of water monomer adsorbed on its surface, contrary to previous findings. For structurally more complex systems of carbonaceous molecules on chemically diverse surfaces, our predictions for adsorption energies are within chemical accuracy of experimental values, demonstrating a good versatility and accuracy of SIE+CCSD(T), paving the way for use in tackling a broad array of material surface problems.
This large scale of quantum chemistry simulations for material surfaces has been enabled by our newly developed SIE+CCSD(T) framework, which leverages the linear scaling nature of fragmentation approaches alongside the computational power of modern GPUs. This efficient framework sets the stage for understanding the broader mechanisms of molecular surface chemistry. Integrating SIE with solvers that can assess excited states will allow for the accurate modeling of photochemistry at surfaces and non-radiative transitions in surface defects and quantum dots. Furthermore, combining this framework with reaction path searching tools will enable detailed mapping of surface reaction processes, and potentially offers predictive understanding of complex chemical reactions on surfaces.
4 Methods
4.1 Framework of SIE+CCSD(T)
Quantum embedding methods strike a balance between accuracy and computational simplicity by selecting a set of overlapping clusters from the system for a high-level numerical method treatment (such as FCI, CCSD(T), QMC, etc), integrated into a low-level (e.g. mean-field) representation of the surrounding environment. These clusters are comprised of local atomic-like fragment spaces over each atomic species, along with a bath space to span the physics of quantum fluctuations from the fragment into its environment as described by the low-level theory. However, the intrinsic mismatch between low-level and high-level solutions in reconstructing the total energy can lead to a lack of clear improvability in quantum embedding making it difficult to systematically improve. SIE provides a more reliable route towards convergence in energetic quantities by expanding the bath space with carefully designed bath natural orbitals (BNOs) relying on a perturbative description of these fluctuations [25].
The SIE framework, outlined in Fig. 1a, requires three key components: a low-level solution, precomputed using mean-field theory (such as the Hartree-Fock method employed in this work); a partitioning strategy designed to generate clusters for quantum embedding; and a BNO threshold to decide the number of BNOs at each level of theory in which the clusters are described. The SIE embedding procedure begins by forming a mean-field bath through the use of the Schmidt decomposition, as employed in density matrix embedding theory (DMET) [27, 28]. The BNOs augment this bath space based on the entanglement between the impurity and its surrounding environment obtained through second-order Møller-Plesset perturbation theory (MP2) over the full system orbitals. Only BNOs entanglement greater than the threshold are included in the cluster at each level of theory. Following this, the many-body problem over each cluster is solved at their respective level of theory, first undergoing MP2, then CCSD on top of the MP2 results, and then the final CCSD(T) calculation. We refer to this process as the SIE+MP2, SIE+CCSD, and SIE+CCSD(T) algorithms, as shown in Fig. 1a. It is important to note that although MP2 and CCSD can be used directly as solvers, CCSD(T) requires directly downfolding correlation into fragments. Therefore, CCSD(T) in SIE is actually a composite approach. A detailed description of perturbative (T) in SIE, along with the energy expression for SIE+CCSD(T), can be found in the SI Sec. LABEL:sec:SIE_(T).
The correlation outside the clusters is estimated by the difference between the full sysetm MP2 and SIE+MP2 [80]. The full system MP2 is not the only candidate to catch outside correlation outside. Any method, which could estimates the correlation in a larger cluster space than SIE+MP2, can do the samething, such as a SIE+MP2 with a lower BNO threshold.
To compute the total energy and other observables, we use the partitioned density matrix approach (PWF-DM) introduced in Ref. [26]. This improves the convergence of energies with respect to bath size by including the inter-cluster interactions between different clusters treated with high-level calculation.
Additionally, we have implemented a GPU-accelerated version of this quantum embedding workflow to dramatically increase computational throughput. As shown in Fig. 1b, the observed complexity of the method has been fitted. Due to noticeable inflection points for those methods in larger systems, only the last three data points were selected for fitting. Interestingly, we observed a linear scaling for SIE+CCSD(T), indicating that with effective engineering optimizations, theoretical bottlenecks do not dominate practical computations for system sizes up to 11,423 orbitals in SIE+CCSD(T). For further discussion on the linear scaling of SIE+CCSD(T), we refer to SI Sec. LABEL:sec:SI_LS_SIE. Specific settings for SIE, including the selection of BNOs, partition schemes, choice of basis sets, as well as detailed error estimation and error correction, can be found in Sec. LABEL:sec:SI_H2O_GrapheneLABEL:sec:SI_setting_summary.
4.2 Adsorption Energy
The adsorption energy is defined as
$\Delta E_{\text{ads}}=E({\text{AB}})-E({\text{A}})-E({\text{B}}),$ (1)
where $\Delta E_{\text{ads}}$ represents the adsorption energy, $E({\text{AB}})$ is the total energy of the system A adsorbed on B, and $E({\text{A}})$ and $E({\text{B}})$ are the total energies of A and B in their isolated thermal equilibrium structures, respectively. In practice, the adsorption energy can be decomposed into two components:
$\Delta E_{\text{ads}}=\Delta E_{\text{int}}+\Delta E_{\text{geom}},$ (2)
where $\Delta E_{\text{int}}$ denotes the energy arising from the interaction between A and B in their equilibrium geometries, and $\Delta E_{\text{geom}}$ is the energy resulting from geometry relaxation of the absorbate relative to its isolated equilibrium structure [59]. Generally, $\Delta E_{\text{geom}}$ is very small for surface adsorption and thus be estimated using DFT calculation as a correction to the adsorption energy. For more details, we refer to Sec. LABEL:sec:SI_geom_relax of the SI. The interaction energy, which constitutes the major part of $\Delta E_{\text{ads}}$, requires accurate computation. However, for the method using Gaussian type orbitals (GTO) as basis sets, the basis set superposition error (BSSE) can arise in adsorption energy calculations [81]. A common approach to address the BSSE is counterpoise correction, applied to $\Delta E_{\text{int}}$ using equation
[…]
Acknowledgments
We extend our gratitude to Dr. Hang Li and ByteDance Research for their invaluable support. We thank Dr. Hong-Zhou Ye for his insightful discussions on the project. We thank Dr. Kaido Sillar and Dr. Joachim Sauer for generously sharing the CPO-27-Mg structure. Additionally, we appreciate Dr. Weiluo Ren, Weizhong Fu and Yibo Wu for their suggestions for this article. J.C. is supported by the National Key R&D Program of China under Grant No. 2021YFA1400500 and the National Science Foundation of China under Grant No. 12334003.
[…]
Contribution
Z.H. and D.L. conceived the study; Z.H. implemented the main code with contributions from Z,G; X.W., J.C., D.L. and G.H.B. suggested the experiments; Z.H. performed simulations and analyzed the results; X.W., C.C, J.C., H.Q.P. and G.H.B. performed the chemical or physical analyses; Z.H., D.L., H.Q.P, J.C., X.W. and C.C. performed figure designing; Z.H. and D.L. organized the project; Z.H. C.C. H.Q.P, X.W., G.H.B., J.C. and D.L. wrote the paper.